# Gravity as a topological gauge theory

### v3

@@ 3.2 Characteristic poylynomials for $G=SO(4,2)/\{\pm I\}$ and associated action
$\left\{\begin{array}[]{ll}&\lambda=0\text{ and }\alpha=\dfrac{k}{2}\beta\\ &\left\{\begin{array}[]{ll}sign(k\ell\ell^{\prime})=-,\quad\Lambda_{0}=-\dfrac{6k}{\ell\ell^{\prime}}>0&\Leftrightarrow G=SO(4,1)\\ k=0&\Leftrightarrow G=ISO(3,1)\\ sign(k\ell\ell^{\prime})=+,\quad\Lambda_{0}=-\dfrac{6k}{\ell\ell^{\prime}}<0&\Leftrightarrow G=SO(3,2).\end{array}\right.\end{array}\right.$ (36)
In the case $\alpha\neq\dfrac{k}{2}\beta$ ($\alpha$ is related to the special conformal transformations) it is interesting to notice that the Palatini action and the cosmological constant term in action (16) respectively become the terms proportional to $\varepsilon_{abcd}R^{ab}\wedge\phi^{cd}$ and $\varepsilon_{abcd}\phi^{ab}\wedge\phi^{cd}$ (it is recalled that $\phi$ was defined in eq.(7)) in the action:
$S[\varpi]$ $=\int_{\mathscr{M}}\biggl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}\bigl{(}R^{ab}\wedge R^{cd}+\dfrac{2}{\ell\ell^{\prime}}R^{ab}\wedge\phi^{cd}+\dfrac{1}{(\ell\ell^{\prime})^{2}}\phi^{ab}\wedge\phi^{cd}\bigr{)}$
$\qquad\quad+\dfrac{1}{8\pi^{2}}\Bigl{(}(r+y)\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2}{\ell\ell^{\prime}}R^{ab}\wedge\phi_{ab}\bigr{)}-\dfrac{4r}{\ell\ell^{\prime}}\eta_{ab}\Pi^{a}\wedge\Theta^{b}-2rd\lambda\wedge d\lambda\Bigr{)}\biggr{)}$ (37)
where the fact that $\alpha^{a}{}_{\mu}\beta^{\,\mu}_{~{}\,\,b}\neq\dfrac{k}{2}\delta^{a}_{b}$ turns out to be an obstruction for $\varepsilon_{abcd}\phi^{ab}\wedge\phi^{cd}$ to be a constant volume form.
@@ 4 Conclusion
We have shown that a large class of models for gravitation, including the standard theory of general relativity, can be solely derived from topological considerations. This has been achieved by considering an action for gravitation as a linear combination of the Euler and Pontrjagin numbers of a manifold $\mathscr{M}$ with specific underlying $G/H$ Cartan geometry. Since the action is expressed as a linear combination of characteristic numbers constructed with various parts of the curvature of a Cartan connection, the action is invariant under the choice of such a Cartan connection up to the integral over $\mathscr{M}$ of an exact term.
@@
Two classes of examples have been studied from which Einstein’s equations (with positive or negative bare cosmological constant $\Lambda_{0}$) can be obtained upon making a restrictive choice (no Pontrjagin numbers).
@@
The first example corresponds to the case of a $\mathfrak{g}=\left\{\begin{array}[]{ll}\mathfrak{so}(4,1)\text{ for }\Lambda_{0}>0\\ \mathfrak{so}(3,2)\text{ for }\Lambda_{0}<0\end{array}\right.$-valued Cartan connection $\varpi$ defined on the $H=SO(3,1)$-principal-bundle $\mathscr{P}$. The computed action gives the Holst action together with the Euler, Pontrjagin and Nieh-Yan topological terms as well as a bare cosmological constant term, see (16). The coupling constants of these topological terms are entirely determined by the three parameters entering the Holst action ($G$, $\gamma$, $r$) and the bare cosmological constant $\Lambda_{0}$.
@@
The second example deals with the Möbius geometry. In this case a very peculiar choice of the "secondary" tetrad $\alpha$ within the $\mathfrak{g}=\mathfrak{so}(4,2)$-valued Cartan connection $\varpi$ defined on the $H=CO(4,2)\ltimes\mathbb{R}^{3,1}$-principal bundle $\mathscr{P}$ yields, either the same results as in the previous class of examples if one drops out dilations, or the Einstein’s equations (modified by the Holst and Nieh-Yan terms) with an additional source term (due to the dilational part of the algebra) for the curvature depending on a particular variation of spin density $-\dfrac{2\pi G}{3}\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;kc}$ and on torsion (if one does not assume $T=0$) when coupled to matter.
In these two classes of examples treated in the paper, the action as linear combination of the Euler and Pontrjagin numbers contains the Holst action showing that the Holst action may have a topological origin. This opens the possibility to figure out a topological formulation of LQG.
Acknowledgements

### v5

@@ 3.2.2 The total action
$\{\lambda=0\text{ and }\alpha=\dfrac{k}{2}\beta\}\,\cap\begin{array}[]{ll}\left\{\begin{array}[]{ll}sign(k\ell\ell^{\prime})=-,\quad\Lambda_{0}=-\dfrac{3k}{\ell\ell^{\prime}}>0&\Leftrightarrow G=SO(4,1)\\ k=0&\Leftrightarrow G=ISO(3,1)\\ sign(k\ell\ell^{\prime})=+,\quad\Lambda_{0}=-\dfrac{3k}{\ell\ell^{\prime}}<0&\Leftrightarrow G=SO(3,2).\end{array}\right.\end{array}$ (37)
$8$-dimensional manifold $\widetilde{\mathscr{M}}$ action
In the case $\alpha\neq\dfrac{k}{2}\beta$ corresponding to the $8$-dimensional manifold $\widetilde{\mathscr{M}}$ ($\alpha$ is related to the special conformal transformations) the action (1) becomes:
$S_{G}[\varpi]$ =~$=\int_{\tilde{\mathscr{M}}}\bigl{(}\overbrace{r\det(\mathds{1}+\dfrac{\bar{\Omega}}{2\pi})}^{P(\bar{\Omega})}+y\det(\mathds{1}+\dfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi})\bigr{)}$ (38)
=~$=\int_{\tilde{\mathscr{M}}}\Bigl{(}-\dfrac{r}{128\pi^{2}}\bigl{(}\operatorname{Tr}(\bar{\Omega}^{2})^{2}-2\operatorname{Tr}(\bar{\Omega}^{4})\bigr{)}-\dfrac{y}{128\pi^{2}}\bigl{(}\operatorname{Tr}(\bar{\Omega}_{\mathfrak{h}}^{2})^{2}-2\operatorname{Tr}(\bar{\Omega}_{\mathfrak{h}}^{4})\bigr{)}\Bigr{)}$
Study of action 38 shall be the subject of another article.
4 Topological total action
In this section we discuss the topological feature of the actions constructed according to the combination (8) we started with. In doing so, we have considered [^21] This ought to be considered as a ”deformation” of the standard Chern-Simons theory. the evaluation of some invariant symmetric polynomials $P(\bar{\Omega}_{\mathfrak{h}}^{n})$ on the $\mathfrak{h}$-part of the Cartan curvature $\bar{\Omega}$ which takes values in a symmetric Lie algebra $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$. One can establish the following
Theorem 4.1.
Let $H\subset G$ be Lie groups such that $G/H$ defines a symmetric space with symmetric Lie algebra $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$ and $\mathscr{M}$ be a manifold of dimension $m$. Let $\varpi=\varpi_{\mathfrak{h}}\oplus\varpi_{\mathfrak{m}}$ be a Cartan connection associated to the Cartan geometry modeled on $G/H$ with curvature $\bar{\Omega}=d\varpi+\tfrac{1}{2}[\varpi,\varpi]=\bar{\Omega}_{\mathfrak{h}}\oplus\bar{\Omega}_{\mathfrak{m}}$. Let $\varpi_{t}=t\varpi$ with curvature $\bar{\Omega}_{t}=d\varpi_{t}+\tfrac{1}{2}[\varpi_{t},\varpi_{t}]$.
Regarding the invariant polynomials $P$ used in our construction two cases are in order:
Case 1. Let $P\in{\cal I}^{n}(G)$ be a $G$-invariant polynomial of degree $n$. Restriction of $P$ to $H$ yields $P\in{\cal I}^{n}(H)$. One has
$(a)$ $\ dP(\bar{\Omega}_{\mathfrak{h}}^{n})=-nP([\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{m}}],\bar{\Omega}_{\mathfrak{h}}^{n-1})=n(n-1)P([\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{h}}],\bar{\Omega}_{\mathfrak{h}}^{n-2},\bar{\Omega}_{\mathfrak{m}})$
[…]
Case 2. Let $P\in{\cal I}^{n}(H)$ be an $H$-invariant polynomial of degree $n$. This case encapsulates the previous one if the $G$-invariance is not fully taken into account. One has
$(c)$ $\ dP(\bar{\Omega}_{\mathfrak{h}}^{n})=-nP([\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{m}}],\bar{\Omega}_{\mathfrak{h}}^{n-1})$
$(d)$ $\ \dfrac{d}{dt}P(\bar{\Omega}_{t,\mathfrak{h}}^{n})=ndP(\varpi_{\mathfrak{h}},\bar{\Omega}_{t,\mathfrak{h}}^{n-1})+nP([\varpi_{t,\mathfrak{m}},\varpi_{\mathfrak{m}}],\bar{\Omega}_{t,\mathfrak{h}}^{n-1})+n(n-1)P(\varpi_{\mathfrak{h}},[\bar{\Omega}_{t,\mathfrak{m}},\varpi_{t,\mathfrak{m}}],\bar{\Omega}_{t,\mathfrak{h}}^{n-2}).$
Proof.
The proof relies on the ones used for the Chern-Weil theorem in [4] or for the Chern-Simons theory related to our situation [10, 30]. Let us first compute
$dP(\bar{\Omega}_{\mathfrak{h}}^{n})$ $=nP(d\bar{\Omega}_{\mathfrak{h}},\bar{\Omega}_{\mathfrak{h}}^{n-1})$
$=nP(d\bar{\Omega}_{\mathfrak{h}}+[\varpi_{\mathfrak{h}},\bar{\Omega}_{\mathfrak{h}}]+[\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{m}}],\bar{\Omega}_{\mathfrak{h}}^{n-1})-nP([\varpi_{\mathfrak{h}},\bar{\Omega}_{\mathfrak{h}}],\bar{\Omega}_{\mathfrak{h}}^{n-1})-nP([\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{m}}],\bar{\Omega}_{\mathfrak{h}}^{n-1})$
$=-nP([\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{m}}],\bar{\Omega}_{\mathfrak{h}}^{n-1})$ (1)
In the middle equation, the first term vanishes thanks to the $\mathfrak{h}$-part of the Bianchi identity, namely, $d\bar{\Omega}_{\mathfrak{h}}+[\varpi_{\mathfrak{h}},\bar{\Omega}_{\mathfrak{h}}]+[\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{m}}]=0$, the second term is nothing but $ad_{\varpi_{\mathfrak{h}}}P(\bar{\Omega}_{\mathfrak{h}}^{n})$ which vanishes by $H$-invariance. The third term will be under discussion whether $P$ is $G$-invariant or if solely its $H$-invariance will be considered.
According to the Chern-Simons theory, let $\varpi_{t}=t\varpi$ be an interpolating family of $\mathfrak{g}$-valued connections, $0\leq t\leq 1$, with curvature $\bar{\Omega}_{t}=\bar{\Omega}_{t,\mathfrak{h}}\oplus\bar{\Omega}_{t,\mathfrak{m}}=d\varpi_{t}+\tfrac{1}{2}[\varpi_{t},\varpi_{t}]$ together with the Bianchi identity $d\bar{\Omega}_{t}+[\varpi_{t},\bar{\Omega}_{t}]=0$. One has $\frac{d}{dt}\bar{\Omega}_{t}=d\varpi+[\varpi_{t},\varpi]$. The standard Chern-Simons derivation leads, on the one hand,
$\frac{d}{dt}P(\bar{\Omega}_{t}^{n})=nP(\frac{d}{dt}\bar{\Omega}_{t},\bar{\Omega}_{t}^{n-1})=nP(d\varpi+[\varpi_{t},\varpi],\bar{\Omega}_{t}^{n-1})$ (2)
and on the other hand,
$ndP(\varpi,\bar{\Omega}_{t}^{n-1})=nP(d\varpi,\bar{\Omega}_{t}^{n-1})-n(n-1)P(\varpi,d\bar{\Omega}_{t},\bar{\Omega}_{t}^{n-2}).$ (3)
The use of the Bianchi identity for $\bar{\Omega}_{t}$ in the right hand side, relies on the Ad-invariance of $P$, which in our situation at hand, depends on either the $G$ or the $H$-invariance of $P$.
Case 1. For $P\in{\cal I}^{n}(G)$, the $Ad(G)$-invariance tells us that
$ad_{\varpi_{t}}P(\varpi,\bar{\Omega}_{t}^{n-1})=P([\varpi_{t},\varpi],\bar{\Omega}_{t}^{n-1})-(n-1)P(\varpi,[\varpi_{t},\bar{\Omega}_{t}],\bar{\Omega}_{t}^{n-2})=0$ (4)
and a straightforward computation yields a link with (2)
$n\,dP(\varpi,\bar{\Omega}_{t}^{n-1})$ $=nP(d\varpi,\bar{\Omega}_{t}^{n-1})-n(n-1)P(\varpi,d\bar{\Omega}_{t},\bar{\Omega}_{t}^{n-2})+ad_{\varpi_{t}}P(\varpi,\bar{\Omega}_{t}^{n-1})$
$=nP(d\varpi+[\varpi_{t},\varpi],\bar{\Omega}_{t}^{n-1})-n(n-1)P(\varpi,d\bar{\Omega}_{t}+[\varpi_{t},\bar{\Omega}_{t}],\bar{\Omega}_{t}^{n-2})$
$=nP(d\varpi+[\varpi_{t},\varpi],\bar{\Omega}_{t}^{n-1})=\frac{d}{dt}P(\bar{\Omega}_{t}^{n})$
where the Bianchi identity for $\bar{\Omega}_{t}$ has been used. The transgression formula gives
[…]
Restriction of $P\in{\cal I}^{n}(G)$ to the subgroup $H\subset G$, requires to consider the $\mathfrak{h}$-part of the curvature from the decomposition $\bar{\Omega}=\bar{\Omega}_{\mathfrak{h}}\oplus\bar{\Omega}_{\mathfrak{m}}$ to obtain the polynomial $P(\bar{\Omega}_{\mathfrak{h}}^{n})$. Going back to (4), since $P\in{\cal I}^{n}(G)$, one can use the $ad(\mathfrak{m})$-invariance,
$ad_{\varpi_{\mathfrak{m}}}P(\bar{\Omega}_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{h}}^{n-1})=P([\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{m}}],\bar{\Omega}_{\mathfrak{h}}^{n-1})+(n-1)P(\bar{\Omega}_{\mathfrak{m}},[\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{h}}],\bar{\Omega}_{\mathfrak{h}}^{n-2})=0.$
This latter relation allows to relax some conditions on torsion of the Cartan connection in order to secure (4) to be 00 ($P(\bar{\Omega}_{\mathfrak{h}}^{n})$ is closed) if one wishes to avoid to consider the $[\mathfrak{m},\mathfrak{m}]$ contributions. Indeed, the Bianchi identity $d\bar{\Omega}_{\mathfrak{m}}+[\varpi_{\mathfrak{h}},\bar{\Omega}_{\mathfrak{m}}]+[\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{h}}]=0$ infers that $[\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{h}}]=0$ if the covariant derivative of the torsion $d\bar{\Omega}_{\mathfrak{m}}+[\varpi_{\mathfrak{h}},\bar{\Omega}_{\mathfrak{m}}]=0$. One thus gets
$(i)$ $\ dP(\bar{\Omega}_{\mathfrak{h}}^{n})=n(n-1)P([\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{h}}],\bar{\Omega}_{\mathfrak{h}}^{n-2},\bar{\Omega}_{\mathfrak{m}})$
[…]
this last equation comes from (5) by considering the expansion of $P(\bar{\Omega}^{n})=P((\bar{\Omega}_{\mathfrak{h}}\oplus\bar{\Omega}_{\mathfrak{m}})^{n})$.
Case 2. For $P\in{\cal I}^{n}(H)$, only the $H$-invariance can be taken into account. Accordingly, one must restrict eqs(2) and (3) to the $\bar{\Omega}_{\mathfrak{h}}$ and $\bar{\Omega}_{t,\mathfrak{h}}$ components only and for which only the $\mathfrak{h}$-part of the Bianchi identity, $(D\bar{\Omega})_{\mathfrak{h}}=0$, can be used. Together with the result (4), all in all, this gives
$(i)$ $~{}dP(\bar{\Omega}_{\mathfrak{h}}^{n})=-nP([\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{m}}],\bar{\Omega}_{\mathfrak{h}}^{n-1})$
$(ii)$ $~{}\dfrac{d}{dt}P(\bar{\Omega}_{t,\mathfrak{h}}^{n})=ndP(\varpi_{\mathfrak{h}},\bar{\Omega}_{t,\mathfrak{h}}^{n-1})+nP([\varpi_{t,m},\varpi_{\mathfrak{m}}],\bar{\Omega}_{t,\mathfrak{h}}^{n-1})+n(n-1)P(\varpi_{\mathfrak{h}},[\bar{\Omega}_{t,\mathfrak{m}},\varpi_{t,\mathfrak{m}}],\bar{\Omega}_{t,\mathfrak{h}}^{n-2})$
where the last two terms could be gathered as the result of $n\,ad_{\varpi_{t,m}}P(\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{h}}^{n-1})$, see (4). But since $P\in{\cal I}^{n}(H)$, the invariance of $P$ under the $\mathfrak{m}$-part of the symmetric Lie algebra $\mathfrak{g}$ no longer holds true. ∎
As can be seen in $(a),(b)(c)$ and $(d)$, the vanishing of the $[\mathfrak{m},\mathfrak{m}]$ commutators removes the obstruction to the Chern-Weil theorem in the case of symmetric Lie algebras. Furthermore, as stated in [31, Proposition 4.1] there is an equivalence between a symmetric Lie algebra $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$ with $[\mathfrak{m},\mathfrak{m}]\subset\mathfrak{h}$ and an abelian reductive decomposition $\mathfrak{g}^{\prime}=\mathfrak{h}\oplus\mathfrak{m}^{\prime}$ with $[\mathfrak{m}^{\prime},\mathfrak{m}^{\prime}]=0$ through a mutation map and where $\mathfrak{m}\simeq\mathfrak{m}^{\prime}$ as $\mathfrak{h}$-modules, (i.e. $[\mathfrak{h},\mathfrak{m}]\subset\mathfrak{m}$ and $[\mathfrak{h},\mathfrak{m}^{\prime}]\subset\mathfrak{m}^{\prime}$ and $\mathfrak{m}\simeq\mathfrak{m}^{\prime}$ as vector spaces, see also [11, Lemma 6.4, p.220]). Exploiting this fact, one has
Corollary 4.1.1.
Another way to reach conditions for the Chern-Weil theorem in both cases is to consider the unique (up to isomorphism) [^22] See [11, Lemma 6.4] and [31, Proposition 4.1]. mutation $\mathfrak{g}\rightarrow\mathfrak{g}^{\prime}$ such that $[\mathfrak{m}^{\prime},\mathfrak{m}^{\prime}]=0$.
In the Lorentzian case of Section 2, this mutation corresponds to setting either $k_{1}$ or $k_{2}=0$, thus obtaining the Minkowski geometry.
On the other hand, for the Lorentz×$\times$×Weyl geometry embedded into the mutated Möbius geometry which is a symmetric Lie algebra, the mutation map giving rise (up to isomorphism) to an abelian reductive decomposition could be
$\varphi_{\mathfrak{g}}=\begin{pmatrix}z&\bar{a}&0\\ b&c&a\\ 0&\bar{b}&-z\end{pmatrix}\longrightarrow\varphi_{\mathfrak{g}^{\prime}}=\mu(\varphi_{\mathfrak{g}})=\left\{\begin{array}[]{ll}\begin{pmatrix}z&0&0\\ \gamma_{1}b&c&\gamma_{2}a\\ 0&0&-z\end{pmatrix}\in\mathfrak{g}^{\prime\prime}=\mathfrak{h}\oplus\mathfrak{m}^{\prime}\\ \\ \begin{pmatrix}z&\bar{\gamma}_{2}\bar{a}&0\\ 0&c&0\\ 0&\bar{\gamma}_{1}\bar{b}&-z\end{pmatrix}\in\mathfrak{g}^{\prime}=\mathfrak{h}\oplus\mathfrak{m}^{\prime\prime}\end{array}\right.$ (9)
where $[\mathfrak{m}^{\prime},\mathfrak{m}^{\prime}]=[\mathfrak{m}^{\prime\prime},\mathfrak{m}^{\prime\prime}]=0$ and $\mathfrak{m}\simeq\mathfrak{m}^{\prime}\simeq\mathfrak{m}^{\prime\prime}$ as $\mathfrak{h}$-modules. They correspond respectively to (4) with $\bar{\gamma}_{1}=\bar{\gamma}_{2}=0$ or with $\gamma_{1}=\gamma_{2}=0$.
Therefore, in the Lorentzian case, see (12), or the mutated Weyl geometry (as a subgeometry of the mutated Möbius geometry) through the identification $\alpha=\dfrac{k}{2}\beta$ together with (9), we respectively have, on the one hand,
$\tfrac{1}{2}[\varpi_{\mathfrak{m}},\varpi_{\mathfrak{m}}]=\left.\begin{cases}\dfrac{k}{\ell^{2}}\begin{pmatrix}\beta\wedge\bar{\beta}&0\\ 0&0\end{pmatrix}=-\dfrac{\Lambda_{0}}{3}\begin{pmatrix}\beta\wedge\bar{\beta}&0\\ 0&0\end{pmatrix}\\ \\ \dfrac{k}{\ell\ell^{\prime}}\begin{pmatrix}0&0&0\\ 0&\beta\wedge\bar{\beta}&0\\ 0&0&0\end{pmatrix}=-\dfrac{\Lambda_{0}}{3}\begin{pmatrix}0&0&0\\ 0&\beta\wedge\bar{\beta}&0\\ 0&0&0\end{pmatrix}\end{cases}\right\}=-\dfrac{\Lambda_{0}}{3}\dfrac{1}{2}\beta^{a}\wedge\beta^{b}J_{ab}$
and, on the other hand,
$[\varpi_{\mathfrak{m}},\bar{\Omega}_{\mathfrak{m}}]$ =$=\begin{cases}\dfrac{k}{\ell^{2}}\begin{pmatrix}\beta\wedge\bar{T}-T\wedge\bar{\beta}&0\\ 0&0\end{pmatrix}=-\dfrac{\Lambda_{0}}{3}\begin{pmatrix}\beta\wedge\bar{T}-T\wedge\bar{\beta}&0\\ 0&0\end{pmatrix}\\ \\ \dfrac{k}{\ell\ell^{\prime}}\begin{pmatrix}0&0&0\\ 0&\beta\wedge\bar{T}-T\wedge\bar{\beta}&0\\ 0&0&0\end{pmatrix}=-\dfrac{\Lambda_{0}}{3}\begin{pmatrix}0&0&0\\ 0&\beta\wedge\bar{T}-T\wedge\bar{\beta}&0\\ 0&0&0\end{pmatrix}\end{cases}$
$=-\dfrac{\Lambda_{0}}{3}\dfrac{1}{2}(\beta^{a}\wedge T^{b}-\beta^{b}\wedge T^{a})J_{ab}.$
[…]
For example, taking the Lorentzian geometry of section 2 we can build the total action: q
$S_{top}$ $=v(S_{G}+S_{M})=v(S_{G}+\mathrm{Re}S_{D})$ (10)
where we tuned $v$ such that $v,v(\dfrac{2}{\gamma}-\dfrac{4\Lambda_{0}Gr}{3\pi})$ tend towards 00. In this way, we retrieve (if we assume $\mathscr{M}$ has no boundary or when neglecting boundary terms) the equations of motion of section 2 with $\Lambda_{0}\rightarrow 0$. Interestingly, we can make $\Lambda_{0}\rightarrow 0^{\pm}$ such that we end up with a dS or AdS spacetime. To some extent, this limit seems to correspond to the interpretation as a "regulator" of the bare cosmological constant in the case of AdS spacetime [20].
Remark.
One is still free to set $v$ such that the coupling constants of the Euler, Pontrjagin and Nieh-Yan terms in the action are respectively $-v\dfrac{3}{64\pi G\Lambda_{0}}$, $v(\dfrac{r}{16\pi^{2}}+\dfrac{3}{32\pi G\Lambda_{0}\gamma})$, $\dfrac{v\Lambda_{0}r}{12\pi^{2}}$ and have non-vanishing values. While the other terms that perturb topological invariance asymptotically vanish.
@@ 5 Conclusion
We have shown that a large class of models for gravitation, including the standard theory of general relativity (modulo the topological Euler density associated to $R$ in the action), can be solely derived from invariant polynomials of the curvature. In particular, LQG with non-trivial Barbero-Immirzi parameter can be obtained by adding the Pontryagin class to the Pfaffian of the Lorentz-valued curvature. In general, only the part related to the Pontryagin class $P(\bar{\Omega})$ is invariant under the choice of Cartan connection up to the integral over $\mathscr{M}$ of an exact term according to the Chern-Weil theorem.
The following tables summarize the roles played by the different characteristic classes and invariant polynomials involved in the actions for the Lorentzian and Möbius geometries. These geometries were studied through mutations allowing to study symmetric types of Cartan geometries with group $H$. The parameters entering into the choice of mutation maps can be related to a bare cosmological constant $\Lambda_{0}$.
Mutated Lorentzian Cartan geometry $\bar{\Omega}_{\mathfrak{h}}=R+\dfrac{k}{\ell^{2}}\beta\wedge\bar{\beta}$ : $\Lambda_{0}=-\dfrac{3k}{\ell^{2}}$
Characteristic class / Terms Theory
Invariant polynomial
$\operatorname{Pf}(\dfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi})$ Palatini + $\Lambda_{0}$ + Euler of R MM gravity
$\mathdutchcal{P}(\bar{\Omega})$ Nieh-Yan + Pontryagin of R Completion of MM gravity to LQG with topological BI parameter
$\det(\mathds{1}+\dfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi})$ $R^{ab}\beta_{a}\beta_{b}$ + Pontryagin of R Completion of Palatini term to Holst
Mutated Möbius Cartan geometry with $\alpha=\dfrac{k}{2}\beta$, $\bar{\Omega}_{\mathfrak{h}}=F+f$ : $\Lambda_{0}=-\dfrac{3k}{\ell\ell^{\prime}}$
Characteristic class / Terms Theory
Invariant Polynomial
$\operatorname{Pf}(\dfrac{F}{2\pi})$ Palatini + $\Lambda_{0}$ + Euler of R MM gravity
$\mathdutchcal{P}(\bar{\Omega})$   Nieh-Yan + Pontrjagin of R               +$d\lambda\wedge d\lambda$ Completion of MM gravity to LQG    with topological BI parameter       and dilational kinetic term
$\det(\mathds{1}+\dfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi})$ $R^{ab}\beta_{a}\beta_{b}$ + Pontryagin of R Completion of Palatini term to Holst
@@
Two classes of examples have been studied from which Einstein’s equations (with positive or negative bare cosmological constant $\Lambda_{0}$) can be obtained.
@@
The first example corresponds to the case of a $\mathfrak{g}=\left\{\begin{array}[]{ll}\mathfrak{so}(4,1)\text{ for }\Lambda_{0}>0\\ \mathfrak{so}(3,2)\text{ for }\Lambda_{0}<0\end{array}\right.$-valued Cartan connection $\varpi$ defined on the $H=SO(3,1)$-principal bundle $\mathscr{P}$. The computed action gives the Holst action together with the Euler, Pontrjagin, Nieh-Yan topological terms as well as a bare cosmological constant term, see (15). The coupling constants of these topological terms are entirely determined by the four parameters entering the Holst action ($G$, $\gamma$), the Nieh-Yan term ($r$) and the bare cosmological constant $\Lambda_{0}$.
@@
The second example deals with the Möbius geometry. In this case restricting the geometry to a $4$-dimensional submanifold $\mathscr{M}$ yields the same results as in the previous class of examples if one drops out dilations. Adding to the original action (3.2) an interaction $T^{a}\wedge\lambda\wedge\beta_{a}$ between dilations and torsion instead leads to Einstein’s equations (modified by the Holst term) with an additional source term (due to the dilational part of the algebra) for the curvature depending on a particular variation of spin density $-\dfrac{2\pi G}{3}\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;kc}$ and on torsion (if one does not assume $T=0$) when coupled to matter.
In these two classes of examples treated in the paper, the Pontrjagin number $\mathdutchcal{P}(\bar{\Omega})$ brings about a topological contribution to the Barbero-Immirzi parameter via the Nieh-Yan density, it also adds the other parity violating term (see [24, 18]) corresponding to the Pontrjagin density of the spacetime curvature $R$.
Finally, the last section addressed the issue of the topological character of the actions. It has been shown that one can recover an action asymptotically invariant under the choice of gauge fields with compact supports and perturbative expansion in the bare cosmological constant $\Lambda_{0}$.
In light of some recent developments in cosmology [32, 33] regarding the Hubble tension, the next step (explored in [34]) shall be to consider spacetime dependent mutations, thus leading to a dynamical cosmological "constant".
Acknowledgments